A 2.50 kg rock is dropped over the edge of a cliff, 35 m above the surface of a river. What is the
speed of the rock just before it strikes the surface of the river?
1
Expert's answer
2020-12-23T07:33:49-0500
"v=\sqrt{2gh}\\v=\sqrt{2(9.8)(35)}=26.2\frac{m}{s}"